Assume that when adults with smartphones are randomly​ selected, 43​% use them in meetings or classes. If 6 adult smartphone users are randomly​ selected, find the probability that at least 3 of them use their smartphones in meetings or classes.
Answer #1

Given data are sample n=6

P= 43% adults use smartphone in meetings or classes = 0.43

Then q=1-p=1-0.43 = 0.57

X= atleast 3 of them use smartphone in meetings or classes

In a binomial distribution p(x)= ncx

P ( atleast 3)= p (x 3) = 1- p (x <3)

= 1- [p (x=0) + p (x=1)+p (x=2)]

=1- [( ) + ()+()]

=1- [0.0343 + 0.1552 +0.2928]

= 1- 0.4823

=0.5177